ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 1991 MS Classification: 34, 81 Mishev, Yordan P., RIMS, Japan On the Fay identity for Kdv tau functions and the identity for the Wronskian of squared solutions of Sturm-Liouville equation We show that the well known identity for the Wronskian of squared solutions of a Sturm-Liouville equation follows from the Fay identity. We also study some odd-order ( (2^{n} − 1)-order , n = 2, 3, …) identities which are specific for tau functions, related to the KdV hierarchy.